# Use Model Averaging instead of Model Selection in Pulsar Timing

### v1

@@ 1 Introduction
Four pulsar timing array (PTA) collaborations recently published evidence for a correlated low-frequency signal thought to arise from a stochastic background of gravitational waves (GWB Agazie et al., 2023c; Reardon et al., 2023a; Antoniadis et al., 2023e; Xu et al., 2023). Such a signal is expected to be generated by a population of supermassive black-hole binaries at the centers of galaxies (Agazie et al., 2023b, a; Antoniadis et al., 2023d, c), but more speculatively also by other sources (Afzal et al., 2023). Bayesian inference lies at the heart of most of these analyses in the literature (van Haasteren et al., 2009; van Haasteren & Vallisneri, 2014), and signal or noise parameters are typically quoted as Bayesian credible intervals. Even Frequentist detection statistics for gravitational waves (GWs Allen & Romano, 2023; Gersbach et al., 2024) usually make use of the noise models obtained by Bayesian inference (Agazie et al., 2023c).
@@ 2.1 Circular analysis in pulsar timing
Model selection for a subsequent analysis on the same data means that in Equation (5) the prior odds $P(M_{i})$ are replaced by something else. Instead of having equal prior odds $P(M_{i})=P(M_{j})$ as assumed in standard single pulsar noise analysis, one sets all $P(M_{j})=0$ except for the one where $j\neq i$, where $P(M_{i}|d)$ had the largest FML during the single pulsar noise analysis. This is a circular use of data, where the results of an analysis are used to inform/tune the model for an analysis of the same data; a well-known practice in the statistics literature called “circular analysis” (Kriegeskorte et al., 2009). Recent examples of this approach are seen in the analysis of the European PTA (EPTA) third data release (Antoniadis et al., 2023a) and the Chinese PTA (CPTA) first data release (Xu et al., 2023), where Bayes Factors are used to deselect noise processes in their GWB analysis (Antoniadis et al., 2023b, e). The practice has been more widespread in the past several years, as the PTA community needed to prune the models for computational efficiency.
@@ 2.3 Correct approach: model averaging
In the literature, model averaging has been applied to pulsar timing using in various ways. General examples are Reversible-Jump Markov MCMC (RJMCMC) methods (Ellis & Cornish, 2016), which are complex and computationally expensive. Product space sampling is more frequently used (Arzoumanian et al., 2018; Taylor et al., 2020) in the PTA community. In the appendix we prove that the usual Log-Uniform priors are equivalent to Spike and Slab priors if one appropriately re-interprets the prior bounds. Spike and Slab priors are, in turn, equivalent to model averaging (Green, 1995). This makes model averaging even more straightforward than RJMCMC or product space sampling. In other words, instead of model selection, the PTA community needs to adopt a workflow where all potential noise sources are included with appropriate priors and so that model averaging is automatically performed. Fortunately, all PTA analyses are already set up to incorporate this workflow. Albeit somewhat slower, the analyses of PTA datasets actually becomes more straightforward when doing model averaging this way. The general advice from the LISA global fit applies here: “Model everything and let the data sort it out” (Cornish & Littenberg, 2015).
@@ 3.2 Procedure and results
The data was generated using the following parameters: $\log_{10}\sigma_{w,a}^{2}=1.5$, $\log_{10}\sigma_{r,a}^{2}=1.5\pm 0.3$ (Gaussian distributed around $1.5$, varied per pulsar), and $\log_{10}h^{2}=1.35$. The priors for signal $\log_{10}h^{2}$ and noise $\log_{10}\sigma_{r,a}$ were set to Log-Uniform priors between $10^{-3}$ and $10^{5}$. Pulsars were positioned uniformly across the sky.
@@
Figure 1: Cross power using the binned pair-covariant correlation statistic (PCOS) of Allen & Romano (2023) for an array of pulsars. The PCOS requires all model parameters as fixed input parameters. For blue “Injected” the true injected parameter values were used. The red “NMOS” has been made with the Full model posterior samples. For orange “MAP Full” the joint Maximum A Posteriori values were used. For green “MAP Data”, the circular analysis Bayesian MAP values were used, where for pulsars with insignificantly detected IRN the noise parameter values were set to zero. It is visible that the MAP Data model has significantly smaller uncertainties than the other models.
Figure 2: Posterior distribution of $\log_{10}h$ for the various models. The blue “HBM” model represents a Hierarchical Bayesian Model (HBM) as in van Haasteren (2024). The orange “Full” model includes all noise processes. The green “Data” model has the noise processes that did not have a Bayes Factor that surpassed the threshold excluded from the model. Although we expect the Full model to show posterior bias due to the lack of an HBM, we can see that the Data model showcases an extra over-estimate of $\log_{10}h$, which makes the posterior distribution incompatible with the injection value.
@@ 4 Circular analysis in PTA literature
In the most-recently released GW searches with PTAs, the Chinese PTA (Xu et al., 2023) and the European PTA (Antoniadis et al., 2023e) prune their models incorrectly using circular analysis on noise processes that are covariant with the GWB. The Parkes PTA (PPTA) does use Bayes Factors for model selection, but the important IRN and dispersion measure variations (DMV) processes that are covariant with the GWB are conservatively all included in the model (Reardon et al., 2023a), which makes the analysis correct. Moreover, the PPTA has done a very exhaustive search for noise processes that the other PTAs have not even attempted, making the analysis quite robust (Reardon et al., 2023b). NANOGrav includes fewer noise components than the PPTA, but IRN and DMVs are included so there are no concerns regarding circular analysis (Agazie et al., 2023d, c).
@@ 5 Conclusions
Bayes Factors are commonly used in PTA projects to select the noise sources that need to be included in the search for GWs as a means to prune the model and to gain computational efficiency. This workflow is statistically incorrect, and can lead to artificially increased GW amplitudes and detection significance reports (Xu et al., 2023; Antoniadis et al., 2023e). Instead, all model components for signal and noise need to be included in the full GW search in PTAs, with Log-Uniform priors on the amplitudes. We have included the proof that this is mathematically equivalent to model averaging over the whole set of possible models. Log-Uniform priors need to be re-interpreted in terms of the prior odds. Exceptions to this rule are signals that we have more prior information on, model components like white noise that are not covariant with signals of interest, and model components that should be modeled with Hierarchical Bayesian Models (van Haasteren, 2024).

### v2

@@ 1 Introduction
Four pulsar timing array (PTA) collaborations recently published evidence for a correlated low-frequency signal thought to arise from a stochastic background of gravitational waves (GWB Agazie et al., 2023a; Reardon et al., 2023a; Antoniadis et al., 2023d; Xu et al., 2023). Such a signal is expected to be generated by a population of supermassive black-hole binaries at the centers of galaxies (Agazie et al., 2023c, d; Antoniadis et al., 2023b, 2024), but more speculatively also by other sources (Afzal et al., 2023). Bayesian inference lies at the heart of most of these analyses in the literature (van Haasteren et al., 2009; van Haasteren & Vallisneri, 2014), and signal or noise parameters are typically quoted as Bayesian credible intervals. Even Frequentist detection statistics for gravitational waves (GWs Allen & Romano, 2023; Gersbach et al., 2024) usually make use of the noise models obtained by Bayesian inference (Agazie et al., 2023a).
@@ 2.1 Circular analysis in pulsar timing
Model selection for a subsequent analysis on the same data means that in Equation (5) the prior odds $P(M_{i})$ are replaced by something else. Instead of having equal prior odds $P(M_{i})=P(M_{j})$ as assumed in standard single pulsar noise analysis, one sets all $P(M_{j})=0$ except for the one where $j\neq i$, where $P(M_{i}|d)$ had the largest FML during the single pulsar noise analysis. This is a circular use of data, where the results of an analysis are used to inform/tune the model for an analysis of the same data; a well-known practice in the statistics literature called “circular analysis” (Kriegeskorte et al., 2009). Recent examples of this approach are seen in the analysis of the European PTA (EPTA) third data release (Antoniadis et al., 2023a) and the Chinese PTA (CPTA) first data release (Xu et al., 2023), where Bayes Factors are used to deselect noise processes in their GWB analysis (Antoniadis et al., 2023c, d). The practice has been more widespread in the past several years, as the PTA community needed to prune the models for computational efficiency.
Using single pulsar noise analysis results to prune the model is not allowed because it effectively restricts the prior range based on the data. However, it is possible that it is discovered during the single pulsar runs that the model needs to be expanded with extra model components that were initially left out. That is appropriate. An example is the extension of the pulsar binary model to include a Shapiro delay component, which are only present for certain orbit orientations.
@@ 2.3 Correct approach: model averaging
In the literature, model averaging has been applied to pulsar timing using in various ways. General examples are Reversible-Jump Markov MCMC (RJMCMC) methods (Ellis & Cornish, 2016), which can be complex to implement efficiently. Product space sampling is more frequently used (Arzoumanian et al., 2018; Taylor et al., 2020) in the PTA community. In the appendix we prove that the usual Log-Uniform priors are equivalent to Spike and Slab priors if one appropriately re-interprets the prior bounds. Spike and Slab priors are, in turn, equivalent to model averaging (Green, 1995). This makes model averaging even more straightforward than RJMCMC or product space sampling. In other words, instead of model selection, the PTA community needs to adopt a workflow where all potential noise sources are included with appropriate priors and so that model averaging is automatically performed. Fortunately, all PTA analyses are already set up to incorporate this workflow. Albeit somewhat slower, the analyses of PTA datasets actually becomes more straightforward when doing model averaging this way. The general advice from the LISA global fit applies here: “Model everything and let the data sort it out” (Cornish & Littenberg, 2015).
@@ 3.2 Procedure and results
The data was generated using the following parameters: $\log_{10}\sigma_{w,a}^{2}=1.5$, $\log_{10}\sigma_{r,a}^{2}=1.5\pm 0.2$ (Gaussian distributed around $1.5$, varied per pulsar), and $\log_{10}h^{2}=1.35$. The priors for signal $\log_{10}h^{2}$ and noise $\log_{10}\sigma_{r,a}$ were set to Log-Uniform priors between $10^{-3}$ and $10^{5}$. Pulsars were positioned uniformly across the sky.
@@
Figure 1: Cross power using the binned pair-covariant correlation statistic (PCOS) of Allen & Romano (2023) for an array of pulsars. The PCOS requires all model parameters as fixed input parameters. For blue “Injected” (leftmost (|)(|...)( | … )) the true injected parameter values were used. The red “NMOS” (second from left (.|..)(.|..)( . | . . )) has been made with the Full model posterior samples. For orange “MAP Full” (second from right (..|.)(..|.)( . . | . )) the joint Maximum A Posteriori values were used. For green “MAP Data” (rightmost (|)(...|)( … | )), the circular analysis Bayesian MAP values were used, where for pulsars with insignificantly detected IRN the noise parameter values were set to zero. It is visible that the MAP Data model has smaller uncertainties than the other models.
Figure 2: Posterior distribution of $\log_{10}h$ for the various models. The blue/solid “HBM” model represents a Hierarchical Bayesian Model (HBM) as in van Haasteren (2024). The orange/dashed “Full” model includes all noise processes. The green/dotted “Data” model has the noise processes that did not have a Bayes Factor that surpassed the threshold excluded from the model. Although we expect the Full model to show posterior bias due to the lack of an HBM, we can see that the Data model showcases an extra over-estimate of $\log_{10}h$, which makes the posterior distribution incompatible with the injection value.
@@ 4 Circular analysis in PTA literature
In the most-recently released GW searches with PTAs, the Chinese PTA (Xu et al., 2023) and the European PTA (Antoniadis et al., 2023d) prune their models incorrectly using circular analysis on noise processes that are covariant with the GWB. The Parkes PTA (PPTA) does use Bayes Factors for model selection, but the important IRN and dispersion measure variations (DMV) processes that are covariant with the GWB are conservatively all included in the model (Reardon et al., 2023a), which makes the analysis correct. Moreover, the PPTA has done a very exhaustive search for noise processes that the other PTAs have not even attempted, making the analysis quite robust (Reardon et al., 2023b). NANOGrav includes fewer noise components than the PPTA, but IRN and DMVs are included so there are no concerns regarding circular analysis (Agazie et al., 2023b, a).
@@ 5 Conclusions
Bayes Factors are commonly used in PTA projects to select the noise sources that need to be included in the search for GWs as a means to prune the model and to gain computational efficiency. This workflow is statistically incorrect, and can lead to artificially increased GW amplitudes and detection significance reports (Xu et al., 2023; Antoniadis et al., 2023d). Instead, all model components for signal and noise need to be included in the full GW search in PTAs, with Log-Uniform priors on the amplitudes. We have included the proof that this is mathematically equivalent to model averaging over the whole set of possible models. Log-Uniform priors need to be re-interpreted in terms of the prior odds. Exceptions to this rule are signals that we have more prior information on, model components like white noise that are not covariant with signals of interest, and model components that should be modeled with Hierarchical Bayesian Models (van Haasteren, 2024).
@@ Acknowledgments
Data availability
The data, code, and figures used in this article can be re-generated with a fully self-contained python script that is available online: https://github.com/vhaasteren/circular_analysis.